OpenStudy (goldrush18):

the table below shows how the total utility that Zach derives from eating candy bars changes as he consumes more and more candy bars each day: Candy Bars Total Utility Marginal Utility 0 - 1 20 2 35 3 12 4 47 5 45 a. Compute the marginal and total utility in the table b. Do the numbers in the table support the law of diminishing marginal utility?explain your answer c. At what point does diminishing marginal utility sets in? d. At what point is utility maximized?

OpenStudy (anonymous):

Total Utility (n+1)= Total Utility (n) + Marginal Utility (n+1 from n) # Total Marginal 1 20 20-0=20 2 35 35-20=15 3 35+12=47 12 4 47 47-47=0 5 45 45-47=-2 Yes law of diminishing marginal utility is supported by the data. At the 5th Candy Bar, Marginal Utility starts diminishing At # of candy bars=3,4 Maximum utility=47 is reached